# Discrete Approximate Circle Bundles

### v2

@@ 3.3 Relationship To True Circle Bundles
Suppose $\{f_{j}\}\in T_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{\sqrt{2}}{28}<\frac{\sqrt{2}}{20}$. By Proposition 3.4, any witness satisfies $\Omega\in\check{Z}_{(3\sqrt{2})\alpha}(\mathcal{U};\underline{O(2)})$, and Proposition 3.18 implies $d_{Z}(\Omega,\widetilde{\Omega})<(27\sqrt{2})\alpha$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. The result then follows from Proposition 3.25.
@@ 4 Effective Computation Of Characteristic Classes
In this section, we give explicit algorithms for computing the characteristic classes of a circle bundle from a discrete approximate cocycle representative. One can then use Construction 3.5 to compute classes from discrete approximate trivializations. We also introduce a notion of persistence for these classes.
We first note some simplifications which occur in the computation of the twisted Euler class when the transition functions are $O(2)$-valued (rather than $\text{Homeo}(\mathbb{S}^{1})$-valued). Let $r:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ denote the section of (2) described in Section 2.3, and let $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ be the group isomorphism described there given by $\varphi(\Lambda,\omega)=\Lambda r(\omega)$. Relative to the reduced structure group $O(2)$, we have $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Moreover, Proposition A.9 implies $\text{Adj}_{r(\omega)}\Lambda=\Lambda^{r(\omega)}$ for any $\Lambda\in O(2)$ and $\omega=\pm 1\in\mathbb{Z}_{2}$, so the group multiplication law in $\mathbb{S}^{1}\rtimes_{r}\mathbb{Z}_{2}$ simplifies to
@@
Thus, the cocycle condition for $(\mathbb{S}^{1}\rtimes\mathbb{Z}_{2})$-valued cocycles is given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
@@
It follows that in this context, the connecting homomorphism $\Delta_{\omega}$ described in Proposition 2.10 reduces to
$\Delta_{\omega}([\Omega])_{jkl}=\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}$ (20)
where $\Omega=(\varphi^{-1})_{*}(\Lambda,\Theta)$. For later use, we define $\text{abs}:O(2)\to SO(2)$ by $\text{abs}(\Omega)=\Omega r(\det(\Omega))$ – note $\varphi^{-1}(\Omega)=(\text{abs}(\Omega),\det(\Omega))$. In fact, it is easy to see that if $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\geq 0$ and $\omega=\det_{*}(\Omega)$, then $\text{abs}_{*}(\Omega)\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$.
Below we give an explicit algorithm for computing characteristic classes from a discrete approximate representative. To prove stability, we first prove a generalization of a technical result from [SP23]. Let $1\to F\to G\xrightarrow{\varphi}H\to 1$ be a central extension of Lie groups with H compact and F discrete. Fix a bi-invariant Riemannian metric on $H$ and use it to metrize $H$ and $G$ with the geodesic distance and $F$ with the distance inherited from $G$. Suppose $F$, $G$ and $H$ have left $\Gamma$-actions by isometries for some abelian group $\Gamma$, and that the maps in the central extension are $\Gamma$-equivariant. Let $\mathcal{U}$ be an open cover of a topological space $B$ such that each set and non-empty binary intersection is locally path connected and simply connected. Let $\text{sys}(H)$ denote the systole of $H$.
@@
For $\varepsilon\leq\frac{\text{sys}(H)}{8}$ and any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\Gamma})$, we have a well-defined map $\Delta_{\omega}:\check{Z}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ given by $\Delta_{\omega}(\Lambda)=[\langle\delta^{1}_{\omega}\Theta\rangle]$, where $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{G,\omega})$ is any lift of $\Lambda$ respect to $\varphi_{*}$. Furthermore, $\Delta_{\omega}$ descends to a map $\Delta_{\omega}:\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ which is stable in the sense that if $\Lambda],[\Lambda^{\prime}]\in\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$ satisfy $d_{H}([\Lambda],[\Lambda^{\prime}])<\frac{\text{sys}(H)}{8}$, then $\Delta_{\omega}([\Lambda])=\Delta_{\omega}([\Lambda^{\prime}])$.
@@
The proof is essentially the same as that of Theorem 6.6 in [SP23]. The fundamental difference is that we are doing cohomology with local coefficients, so the coboundary maps are slightly different. On the other hand, because $\Gamma$ acts by isometries, the same arguments apply. In particular, for any $\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$, one can use the same steps shown in [SP23] to verify the 2-cochain given by $\beta_{jkl}=\langle(\delta_{\omega}^{1}\Theta)_{jkl}\rangle$ is a genuine cocycle, and the stability arguments are identical.
@@
For $\varepsilon\leq 1$, we have a well-defined map $\text{Cl}:\check{H}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{O(2)})\to\coprod_{[\omega]\in\check{H}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})}\check{H}^{2}(B;\mathbb{Z}_{\omega})$ given by $\Omega\longmapsto([\det_{*}\Omega],[\Delta_{\omega}(\text{abs}_{*}\Omega)])$. This map is 1-stable in the sense that $\text{Cl}(\Omega)=\text{Cl}(\Omega^{\prime})$ whenever $d_{H}(\Omega,\Omega^{\prime})<1$. Moreover, if $\varepsilon<\frac{1}{9}$, then $\text{Cl}(\Omega)$ returns the first Stiefel-Whitney class and twisted Euler class of the true circle bundle represented by $\Pi_{Z}(\Omega)$.
@@
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\leq 1$. Lemma 1.3 of [SP23] implies $\omega=\det_{*}(\Omega)$ is a cocycle and that $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any other $\Omega^{\prime}\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Omega^{\prime})<2$. If $\varepsilon<1/9$, then by Proposition 3.18, there exists a true cocycle $\Omega^{\prime}$ such that $d_{Z}(\Omega,\Omega^{\prime})<1$, so $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any such $\Omega^{\prime}$. We conclude that the first Stiefel-Whitney class of any such $\Omega^{\prime}$ is the same.
Now, recall central extension (9). Let $SO(2)_{g}$ denote the Lie group $SO(2)$ endowed with the geodesic distance metric (via the natural identification of $SO(2)$ with $\mathbb{S}^{1}$). The induced metrics on $\mathbb{R}$ and $\mathbb{Z}$ are then the usual ones, and all three groups have $\mathbb{Z}_{2}$-actions by isometries such that the inclusion and $\exp$ maps are equivariant. Thus, Proposition 4.1 gives a well-defined map $\Delta_{\omega}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ for all $\varepsilon\leq\frac{sys(SO(2)_{g})}{8}$, where $\omega=\det_{*}(\Omega)$. By Lemma A.18 from [SP23], $\frac{sys(SO(2)_{g})}{8}=\frac{2\pi\sqrt{2}}{8}=\frac{\pi\sqrt{2}}{4}$ with respect to geodesic distance. To obtain a result in terms of the Frobenius norm, we note Corollary A.17 from [SP23] implies we have an inclusion $\check{Z}_{1}^{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})\subset\check{Z}^{1}_{\frac{\pi\sqrt{2}}{4}}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})$. Thus, $\text{abs}_{*}(\Omega)\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$, so $\Delta_{\omega}(\text{abs}_{*}(\Omega))\in\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ (and therefore $\text{Cl}(\Omega)$) is well-defined.
@@
Below is an explicit algorithm for computing characteristic class representatives.
Input: A cochain $\Omega\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ subordinate to a finite open cover $\mathcal{U}$.
Output: A pair of cochains $\omega\in\check{C}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}\in\check{C}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$.
  1. 1.
    Compute $\omega_{jk}=\det(\Omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
  2. 2.
    For each $(jk)\in\mathcal{N}(\mathcal{U})$, compute $\Lambda_{jk}=\Omega_{jk}r(\omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$, where $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Then, choose a lift $\Theta_{jk}\in\mathbb{R}$ of each $\Lambda_{jk}$ via the $\exp$ map defined in (9).
  3. 3.
    Compute $\tilde{e}_{jkl}=\langle\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}\rangle$ for each $(jkl)\in\mathcal{N}(\mathcal{U})$.

### v3

@@ 3.5 Relationship To True Circle Bundles
Suppose $\mathfrak{F}\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for $\alpha<\frac{1}{28}<\frac{\sqrt{2}}{28}<\frac{\sqrt{2}}{20}$. By Proposition 3.15, any witness satisfies $\Omega\in\check{Z}_{(3\sqrt{2})\alpha}\left(\mathcal{U};\underline{O(2)}\right)$, and Proposition 3.29 implies $\delta=d_{Z}(\Omega,\widetilde{\Omega})<(27\sqrt{2})\alpha$, where $\widetilde{\Omega}=\Pi_{Z}(\Omega)$. Thus,
$\alpha+\delta\ <\ (1+27\sqrt{2})\alpha\ <\ \frac{1+27\sqrt{2}}{28}\ <\ \sqrt{2}$
so the result follows by Proposition 3.38.
To describe conditions under which a discrete approximate circle bundle can be uniquely identified with a class of true bundles, we introduce the following notation:
Definition 3.40.
For any $r>0$, $\beta,\varepsilon\geq 0$ and collection of open sets $\mathcal{U}$, define
$\tau(r,\beta,\varepsilon,\mathcal{U})=\frac{1}{r}\left(\sup_{j\in J}\ \text{diam}(U_{j})+10\varepsilon+9\beta\right),\hskip 28.45274pt\tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})=2\sin^{-1}\left(\frac{\tau(r,\beta,\varepsilon,\mathcal{U})}{2}\right)$ (24)
Observe that if $\pi:X\to B$ is a discrete $(r,\beta,\varepsilon)$-approximate circle bundle and $\Phi\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$, then
$\tau(\Phi)\ \leq\ \tau(r,\beta,\varepsilon,\mathcal{U}),\hskip 28.45274pt\widetilde{\tau}(\Phi)\ \leq\ \widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
The following result shows that witnesses to systems of discrete approximate local angle functions emerging from trivializations subordinate to the same cover are roughly related by a change of gauge:
Lemma 3.41.
Let $\pi:X\to B$ be a discrete $(r,\beta,\varepsilon)$-approximate circle bundle for $r>0$, $\beta,\varepsilon\geq 0$ with trivializing cover $\mathcal{U}=\{U_{j}\}_{j\in J}$, and let $\Omega,\Omega^{\prime}\in\check{C}^{1}(\mathcal{U};\underline{O(2)})$ respectively witness that $\mathfrak{F}(\Phi),\mathfrak{F}(\Phi^{\prime})\in\Gamma_{\alpha}(\pi,\mathcal{U})$ for some $\alpha\leq\max\{\alpha(\Phi),\alpha(\Phi^{\prime})\}$. There exists some $\zeta\in\check{C}^{0}(\mathcal{U};\underline{O(2)})$ such that
$\|\Omega_{jk}^{\prime}-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}\|_{F}\quad\leq\quad 5\sqrt{2}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
for all $(jk)\in\mathcal{N}(\mathcal{U})$.
Proof.
Define
$\tau(\Phi,\Phi^{\prime})=\sup_{j\in J}\ \tau(\varphi_{j},\varphi_{j}^{\prime}),\hskip 28.45274pt\tilde{\tau}(\Phi,\Phi^{\prime})=2\sin^{-1}\left(\frac{\tau(\Phi,\Phi^{\prime})}{2}\right)$
and observe
$\tau(\Phi,\Phi^{\prime})\ \leq\ \tau(r,\beta,\varepsilon,\mathcal{U})\hskip 28.45274pt\tilde{\tau}(\Phi,\Phi^{\prime})\ \leq\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
Lemma 3.10 implies that for each $j$, there exists some $\zeta_{j}\in O(2)$ such that $d_{\infty}(f_{j}^{\prime},\zeta_{j}f_{j})\leq\tau(\Phi,\Phi^{\prime})$. Thus, for any $x\in\pi^{-1}(U_{jk})$,
$|f_{j}^{\prime}(x)-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|\quad$ $\leq\quad|f_{j}^{\prime}(x)-\zeta_{j}f_{j}(x)|\ +\ |\zeta_{j}f_{j}(x)-\zeta_{j}\Omega_{jk}f_{k}(x)|\ +\ |\zeta_{j}\Omega_{jk}f_{k}(x)-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|$
$=\quad|f_{j}^{\prime}(x)-\zeta_{j}f_{j}(x)|\ +\ |f_{j}(x)-\Omega_{jk}f_{k}(x)|\ +\ |\zeta_{k}f_{k}(x)-f_{k}^{\prime}(x)|$
$\leq\quad\tilde{\tau}(\Phi,\Phi^{\prime})+\tilde{\tau}(\Phi)+\tilde{\tau}(\Phi^{\prime},\Phi)$
$=\quad 2\tilde{\tau}(\Phi,\Phi^{\prime})+\tilde{\tau}(\Phi)$
which implies
$|\Omega_{jk}^{\prime}f_{k}^{\prime}(x)-\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|\quad$ $\leq\quad|\Omega_{jk}^{\prime}f_{k}^{\prime}(x)-f_{j}^{\prime}(x)|+|f_{j}^{\prime}(x),\zeta_{j}\Omega_{jk}\zeta_{k}^{-1}f_{k}^{\prime}(x)|$
$\leq\quad\tilde{\tau}(\Phi^{\prime})+(2\tilde{\tau}(\Phi,\Phi^{\prime})+\tilde{\tau}(\Phi))$
$\leq\quad 4\tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
for all $x\in\pi^{-1}(U_{jk})$. Using (17) and applying Proposition A.12 to the matrix $\Omega_{jk}^{\prime}-\xi_{j}\Omega_{jk}\xi_{k}^{\prime}$ , we obtain
$\|\Omega_{jk}^{\prime}-\xi_{j}\Omega_{jk}\xi_{k}^{\prime}\|_{F}\quad\leq\quad\frac{4\sqrt{2}}{1-2\sin^{-1}\left(\frac{\varepsilon}{2r}\right)}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\quad\leq\quad 5\sqrt{2}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
We now arrive at the main result of this section:
Theorem 3.42.
Let $\xi=\pi:X\to B$ be a discrete $(r,\beta,\varepsilon)$-approximate circle bundle for some $r,\beta,\varepsilon\geq 0$, and let $\mathcal{U}=\{U_{j}\}_{j\in J}$ be a trivializing cover. If $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\leq\frac{\sqrt{2}}{36}$, then for any $\Phi\in T_{r,\beta,\varepsilon}(\pi,\mathcal{U})$, the cocycle $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi))\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ is well-defined, and its cohomology class does not depend on $\Phi$.
Proof.
By Proposition 3.11, we have $\mathfrak{F}(\Phi)\in\Gamma_{\alpha(\Phi)}(\pi,\mathcal{U})$, where $\alpha(\Phi)$ is defined by (16). Thus, Proposition 3.15 implies $\text{wit}(\mathfrak{F}(\Phi))\in\check{Z}^{1}_{(3\sqrt{2})\alpha(\Phi)}(\mathcal{U};\mathcal{C}_{O(2)})$. By Proposition 3.29, $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi))\in\check{Z}^{1}(\mathcal{U};\mathcal{C}_{O(2)})$ is well-defined as long as $(3\sqrt{2})\alpha(\Phi)\leq\frac{1}{2}$, which is satisfied as long as $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\leq\frac{\sqrt{2}}{15}$.
Now, suppose we have some other $\Phi^{\prime}\in T_{r,\varepsilon,\beta}(\pi,\mathcal{U})$. Let $\Omega=\text{wit}(\mathfrak{F}(\Phi))$ and let $\Omega^{\prime}=\text{wit}(\mathfrak{F}(\Phi^{\prime}))$ Proposition 3.30 implies that if $3\sqrt{2}\max\{\alpha(\Phi),\alpha(\Phi^{\prime})\}+d_{H}([\Omega],[\Omega^{\prime}])\leq\frac{1}{2}$, then $\Pi_{Z}(\Omega^{\prime})]=[\Pi_{Z}(\Omega)]$. By Lemma 3.41, $d_{H}([\Omega],[\Omega^{\prime}])\leq 5\sqrt{2}\ \tilde{\tau}(r,\beta,\varepsilon,\mathcal{U}),$ so
$3\sqrt{2}\max\{\alpha(\Phi),\alpha(\Phi^{\prime})\}+d_{H}([\Omega],[\Omega^{\prime}])\quad\leq\quad\frac{5}{4}(3\sqrt{2})\widetilde{\tau}(\Phi)+(5\sqrt{2})\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\quad\leq\quad(9\sqrt{2})\tilde{\tau}(r,\beta,\varepsilon,\mathcal{U})$
We conclude that $\Pi_{Z}(\Omega)]=[\Pi_{Z}(\Omega^{\prime})]$ as long as $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{U})\leq\frac{1}{18\sqrt{2}}=\frac{\sqrt{2}}{36}$.
Remark 3.43.
The cohomology class assigned to $\xi$ in the proposition above is also independent of the trivializing cover $\mathcal{U}$ following sense: suppose $\Phi^{\prime}\in T_{r,\beta,\varepsilon}(\pi,\mathcal{V})$ for some other trivializing cover $\mathcal{V}=\{V_{j^{\prime}}\}_{j^{\prime}\in J^{\prime}}$ satisfying $\widetilde{\tau}(r,\beta,\varepsilon,\mathcal{V})<\frac{\sqrt{2}}{36}$. If $\mathcal{W}$ is a common refinement of $\mathcal{U}$ and $\mathcal{V}$, then the images of $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi))]$ and $\Pi_{Z}\circ\text{wit}(\mathfrak{F}(\Phi^{\prime}))]$ with respect to the induced inclusions coincide in $\check{H}^{1}(\mathcal{W};\mathcal{C}_{O(2)})$.
@@ 4 Effective Computation Of Characteristic Classes
In this section, we give explicit algorithms for computing the characteristic classes of a circle bundle from a discrete approximate cocycle representative. One can then use Construction 3.16 to compute classes from discrete approximate trivializations. We also introduce a notion of persistence for these classes.
We first note some simplifications which occur in the computation of the twisted Euler class when the transition functions are $O(2)$-valued (rather than $\text{Homeo}(\mathbb{S}^{1})$-valued). Let $r:\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ denote the section of (2) described in Section 2.3, and let $\varphi:\text{Homeo}_{+}(\mathbb{S}^{1})\rtimes_{r}\mathbb{Z}_{2}\to\text{Homeo}(\mathbb{S}^{1})$ be the group isomorphism described there given by $\varphi(\Lambda,\omega)=\Lambda r(\omega)$. Relative to the reduced structure group $O(2)$, we have $r(\pm 1)=\left(\begin{array}[]{cc}1&0\\ 0&\pm 1\\ \end{array}\right)$. Moreover, Proposition A.11 implies $\text{Adj}_{r(\omega)}(\Lambda)=\Lambda^{r(\omega)}$ for any $\Lambda\in O(2)$ and $\omega=\pm 1\in\mathbb{Z}_{2}$, so the group multiplication law in $\mathbb{S}^{1}\rtimes_{r}\mathbb{Z}_{2}$ simplifies to
@@
The cocycle condition for $(\mathbb{S}^{1}\rtimes\mathbb{Z}_{2})$-valued cocycles is therefore given explicitly by $(\Lambda_{jk},\omega_{jk})\cdot(\Lambda_{kl},\omega_{kl})=(\Lambda_{jl},\omega_{jl})$, or, equivalently,
@@
Thus, for any $\Lambda]\in\check{H}^{1}(\mathcal{U};\mathcal{C}_{O(2),\omega})$, the connecting homomorphism $\Delta_{\omega}$ described in Proposition 2.11 returns a class $\Delta_{\omega}([\Lambda])\in\check{H}^{2}(\mathcal{U};\underline{\mathbb{Z}})$ represented by the cocycle whose $(jkl)$-component is given by
$\omega_{jk}\Theta_{kl}(b)-\Theta_{jl}(b)+\Theta_{jk}(b)$ (27)
for all $b\in U_{j}\cap U_{k}\cap U_{l}$, where $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{R}})$ is a lift of $\Lambda$ via $\exp_{*}$.
For later use, we define
$\begin{array}[]{rccl}\text{abs}:&O(2)&\longrightarrow&SO(2)\\ &A&\mapsto&Ar(\det(A))\end{array}$ (28)
and note that $\varphi^{-1}(A)=(\text{abs}(A),\det(A))$. In fact, it is easy to see that if $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\geq 0$ and $\omega=\det_{*}(\Omega)$, then the cochain $\Lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{SO(2)})$ given by $\Lambda_{jk}(b)=\text{abs}(\Omega_{jk}(b))$ for all $U_{j}\cap U_{k}\in\mathcal{N}(\mathcal{U})$ and $b\in U_{j}\cap U_{k}$ lies in $\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$. We will denote the cochain $\Lambda$ so defined by $\text{abs}(\Omega)$.
Given a real number $x$, we use the notation $\langle x\rangle$ to denote the nearest integer to $x$ (if the integer is not unique, we take the larger one). Similarly, for any cochain $\lambda\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{\mathbb{R}})$, we denote the nearest integer-valued cochain (with respect to $d_{Z}$) by $\langle\lambda\rangle$.
Below we give an explicit algorithm for computing characteristic classes from a discrete approximate representative. To prove stability, we first prove a generalization of a technical result from [24] (Theorem 6.6). Let $1\to F\to G\xrightarrow{\psi}H\to 1$ be a central extension of Lie groups with H compact and F discrete. Fix a bi-invariant Riemannian metric on $H$ and use it to metrize $H$ and $G$ with the geodesic distance and $F$ with the distance inherited from $G$. Suppose $F$, $G$ and $H$ have left $\Gamma$-actions by isometries for some abelian group $\Gamma$, and that the maps in the central extension are $\Gamma$-equivariant. Let $\mathcal{U}$ be an open cover of a topological space $B$ such that each set and non-empty binary intersection is locally path connected and simply connected. Let $\text{sys}(H)$ denote the systole of $H$.
@@
For $\varepsilon\leq\frac{\text{sys}(H)}{8}$ and any $\omega\in\check{Z}^{1}(\mathcal{U};\underline{\Gamma})$, we have a well-defined map $\beta_{\omega}:\check{Z}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ given by $\Delta_{\omega}(\Lambda)=[\langle\beta^{1}_{\omega}\Theta\rangle]$, where $\Theta\in\check{C}^{1}(\mathcal{U};\mathcal{C}_{G,\omega})$ is any lift of $\Lambda$ respect to $\psi_{*}$ and $\langle\cdot\rangle$ denotes the unique nearest integer Furthermore, $\Delta_{\omega}$ descends to a map $\Delta_{\omega}:\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})\to\check{H}^{2}(\mathcal{U};F_{\omega})$ which is stable in the sense that if $\Lambda],[\Lambda^{\prime}]\in\check{H}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$ satisfy $d_{H}([\Lambda],[\Lambda^{\prime}])<\frac{\text{sys}(H)}{8}$, then $\Delta_{\omega}([\Lambda])=\Delta_{\omega}([\Lambda^{\prime}])$.
@@
The proof is essentially the same as that of Theorem 6.6 in [24]. The fundamental difference is that we are doing cohomology with local coefficients, so the coboundary maps are slightly different. On the other hand, because $\Gamma$ acts by isometries, the same arguments apply. In particular, for any $\Lambda\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{H,\omega})$, one can use the same steps shown in [24] to verify the 2-cochain given by $\beta_{jkl}=\langle(\delta_{\omega}^{1}\Theta)_{jkl}\rangle$ is a genuine cocycle, and the stability arguments are identical.
@@
For $\varepsilon\leq 1$, we have a well-defined function
$\text{Cl}:\check{H}_{\varepsilon}^{1}(\mathcal{U};\mathcal{C}_{O(2)})\to\coprod_{[\omega]\in\check{H}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})}\check{H}^{2}(B;\mathbb{Z}_{\omega})$
given by $\Omega\longmapsto([\det_{*}\Omega],\Delta_{\omega}(\text{abs}(\Omega)))$. This map is 1-stable in the sense that $\text{Cl}(\Omega)=\text{Cl}(\Omega^{\prime})$ whenever $d_{H}(\Omega,\Omega^{\prime})<1$. Moreover, if $\varepsilon<\frac{1}{9}$, then $\text{Cl}(\Omega)$ returns the first Stiefel-Whitney class and twisted Euler class of the true circle bundle represented by $\Pi_{Z}(\Omega)$.
@@
Suppose $\Omega\in\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{O(2)})$ for some $\varepsilon\leq 1$. Lemma 1.3 of [24] implies $\omega=\det_{*}(\Omega)$ is a cocycle and that $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any other $\Omega^{\prime}\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{O(2)})$ such that $d_{Z}(\Omega,\Omega^{\prime})<2$. If $\varepsilon<1/9$, then by Proposition 3.29, there exists a true cocycle $\Omega^{\prime}$ such that $d_{Z}(\Omega,\Omega^{\prime})<1$, so $\det_{*}(\Omega)=\det_{*}(\Omega^{\prime})$ for any such $\Omega^{\prime}$. We conclude that the first Stiefel-Whitney class of any such $\Omega^{\prime}$ is the same.
Now, recall central extension (8). Let $SO(2)_{g}$ denote the Lie group $SO(2)$ endowed with the geodesic distance metric (via the natural identification of $SO(2)$ with $\mathbb{S}^{1}$). The induced metrics on $\mathbb{R}$ and $\mathbb{Z}$ are then the usual ones, and all three groups have $\mathbb{Z}_{2}$-actions by isometries such that the inclusion and $\exp$ maps are equivariant. Thus, Proposition 4.1 gives a well-defined map $\Delta_{\omega}:\check{Z}^{1}_{\varepsilon}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})\to\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ for all $\varepsilon\leq\frac{sys(SO(2)_{g})}{8}$, where $\omega=\det_{*}(\Omega)$. By Lemma A.18 from [24], $\frac{sys(SO(2)_{g})}{8}=\frac{2\pi\sqrt{2}}{8}=\frac{\pi\sqrt{2}}{4}$ with respect to geodesic distance. To obtain a result in terms of the Frobenius norm, we note Corollary A.17 from [24] implies we have an inclusion $\check{Z}_{1}^{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})\subset\check{Z}^{1}_{\frac{\pi\sqrt{2}}{4}}(\mathcal{U};\mathcal{C}_{SO(2)_{g},\omega})$. Thus, $\text{abs}(\Omega)\in\check{Z}^{1}_{1}(\mathcal{U};\mathcal{C}_{SO(2),\omega})$, so $\Delta_{\omega}(\text{abs}(\Omega))\in\check{H}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$ (and therefore $\text{Cl}(\Omega)$) is well-defined.
@@
Below is an explicit algorithm for computing characteristic class representatives:
Algorithm 1 (Computation Of Characteristic Classes).
1:
2:A cochain $\Omega\in\check{C}^{1}\!\left(\mathcal{U};\underline{O(2)}\right)$ subordinate to a good, finite open cover $\mathcal{U}$.
3:A pair of cochains $\omega\in\check{C}^{1}(\mathcal{U};\underline{\mathbb{Z}_{2}})$ and $\tilde{e}\in\check{C}^{2}(\mathcal{U};\mathbb{Z}_{\omega})$
4:Compute $\omega_{jk}=\det(\Omega_{jk})$ for all $(jk)\in\mathcal{N}(\mathcal{U})$.
5:For each $(jk)\in\mathcal{N}(\mathcal{U})$, set $\Lambda_{jk}=\Omega_{jk}\,r(\omega_{jk})$ with $r(\pm 1)=\begin{bmatrix}1&0\\[2.0pt] 0&\pm 1\end{bmatrix}$, then choose a lift $\Theta_{jk}\in\mathbb{R}$ via the $\exp$ map in (8).
6:For each $(jkl)\in\mathcal{N}(\mathcal{U})$, compute $\tilde{e}_{jkl}=\big\langle\,\omega_{jk}\Theta_{kl}-\Theta_{jl}+\Theta_{jk}\,\big\rangle$.